OpenStudy (anonymous):

calc help, product rule xe^(x^(2)) f'*g + g'*f e^(x^(2)) + ? * x not sure what the derivative of e^(x^(2)) is

OpenStudy (ash2326):

Product rule \[d(f\times g)=f'\times g+f\times g'\]

OpenStudy (anonymous):

if the x didnt have a ^2 id be fine but i just dont how to solve for that

OpenStudy (klimenkov):

Use \[\frac{d}{dx}e^{f(x)}=e^{f(x)}\cdot f'(x)\]

OpenStudy (anonymous):

\[ (e^{x^2})' =e^{x^2} \cdot 2x \]

OpenStudy (ash2326):

here we have \[(xe^{x^2})=e^{x^2}\frac d {dx} (x)+x\frac d {dx} (e^{x^2})\times {\frac{d}{dx}x^2}\]

OpenStudy (anonymous):

so then the answer of the problem would be 1 + 2x^2?

OpenStudy (ash2326):

multiplied by \(e^{x^2}\)

OpenStudy (anonymous):

ok thanks alot

OpenStudy (ash2326):

@kingpin do you understand this?

OpenStudy (anonymous):

mhmm